# Cutoff for Contingency Table Random Walks

### v1

@@ Abstract
We study the Diaconis-Gangolli random walk on $n\times n$ contingency tables and its analog on $1\times n$ contingency tables, both over ${\mathbb{Z}}/q{\mathbb{Z}}$. In the $1\times n$ case, we prove that the random walk exhibits cutoff at time $\dfrac{nq^{2}\log(n)}{8\pi^{2}}$ when $q\gg n$; in the $n\times n$ case, we establish cutoff for the random walk at time $\dfrac{n^{2}q^{2}\log(n)}{8\pi^{2}}$ when $q\gg n^{2}$. We also show that a general class of random walks on $({\mathbb{Z}}/q{\mathbb{Z}})^{n}$ with a marginal incremental variance $\dfrac{\sigma^{2}}{n}$ (when mapped to $({\mathbb{Z}}\cap[-q/2,q/2))^{n}$) has cutoff at time $\dfrac{nq^{2}\log(n)}{4\pi^{2}\sigma^{2}}$.
@@ 1.1 Motivation and Basic Definitions
A contingency table is a matrix of integer entries with prescribed row and column sums. Enumerating the contingency tables with given marginals is closely related to counting a variety of other combinatorial objects, including permutations with descent restrictions, double cosets of the symmetric groups, certain quantities about the induced representations of the symmetric groups, and some special types of Young tableaux. The name “contingency table” is commonly used in the context of statistics, and is linked to Fisher’s exact test of association and goodness of fit, along with the approximate $\chi^{2}$ test. Randomly sampling from the set of contingency tables with fixed margins can be used to form an approximate p-value for the test of association when the $\chi^{2}$ approximation is not well met.
Markov chain Monte Carlo methods can be used to randomly sample and approximately count contingency tables with fixed row and column sums. One of the famous MCMC algorithms was described in [7] and [11], which gives rise to the Diaconis-Gangolli random walk on contingency tables over ${\mathbb{Z}}_{q}={\mathbb{Z}}/q{\mathbb{Z}}$. Nestoridi and Nguyen [24] established cutoff for the discrete-time Diaconis-Gangolli random walk on $n\times n$ contingency tables over ${\mathbb{Z}}_{q}$, as $n\to\infty$. We show cutoff in continuous time when $q$ is sufficiently large relative to $n^{2}$, while Nestoridi and Nguyen needed $\log(q)=o\left(\dfrac{\sqrt{\log(n)}}{\log\log n}\right)$ to achieve cutoff.
@@
In addition, we generalize our methods and establish cutoff for a class of random walks on $\mathbb{Z}_{q}^{n}$ satisfying assumptions on the correlations between the coordinates of a single increment. This expands on the results of finding cutoff for non-local random walks on $\mathbb{Z}_{q}^{n}$. Previously, Nestoridi [23] found mixing times and regimes of cutoff for a non-local random walk on ${\mathbb{Z}}_{2}^{n}=({\mathbb{Z}}/2{\mathbb{Z}})^{n}$ such that $k$ coordinates are randomly selected and flipped in a single iteration. In contrast, our results show cutoff for random walks on ${\mathbb{Z}}_{q}^{n}$ with non-nearest-neighbor steps when both $n\to\infty$ and $q=q(n)\to\infty$.
@@
We now define the dynamics of the Diaconis-Gangolli random walk and note that we will study this model in continuous time. We start with the $1\times n$ contingency tables.
@@
    Flip a fair coin. If heads, increment the first coordinate selected by $+1$ and increment the second coordinate selected by $-1$; if tails, increment the first coordinate selected by $-1$ and increment the second coordinate selected by $+1$.
The analogous definition and dynamics for $n\times n$ contingency tables are the following.
@@
We will take $m=n$ and consider the rate-1 continuous-time Diaconis-Gangolli random walk on $n\times n$ contingency tables over ${\mathbb{Z}}_{q}$, each of whose jumps is governed by the following rule:
@@
    Flip a fair coin, and consider the $n\times n$ matrix $T=T(i,i^{\prime},j,j^{\prime})$ whose $(i,j)$-th and $(i^{\prime},j^{\prime})$-th entries are $+1$, $(i,j^{\prime})$-th and $(i^{\prime},j)$-th entries are $-1$, and zero elsewhere,
@@
    If heads, add $T$ to the current matrix; if tails, add $-T$ to the current matrix.
@@
for $\varepsilon\in(0,1)$. In particular, we show that both the walks on $1\times n$ and $n\times n$ contingency tables exhibit cutoff as $n\to\infty$. We say that a sequence of Markov chains indexed by $n=1,2,\ldots$ has cutoff if for all $\varepsilon\in(0,1)$,
@@
In other words, for large enough $n$, there will be a sharp transition for $d(t)$ from almost $1$ to almost 00.
@@ 1.2 Notations
Throughout the paper, we will use ${\mathbb{Z}}_{q}$ to denote the cyclic group ${\mathbb{Z}}/q{\mathbb{Z}}$ of $q$ elements. We do not need $q$ to be a prime number, but we intrinsically require $q=q(n)$ to increase along with $n$.
@@
  • $f={\mathcal{O}}(g)$, or equivalently, $f\lesssim g$, if lim $\limsup f/g<\infty$, and
@@
All the limits are taken as $n\to\infty$ unless otherwise specified. Lastly, we will write =${\mathbb{N}}=\{0,1,2,3,\dots\}$ for the set of nonnegative integers.
@@ 1.3 Main Results
Let $\varepsilon\in(0,1)$ be arbitrary. For the rate-$1$ continuous-time Diaconis-Gangolli random walk on $1\times n$ contingency tables, we have the following.
@@
Let $\varepsilon\in(0,1)$ be arbitrary. For the rate-$1$ continuous-time Diaconis-Gangolli random walk on $n\times n$ contingency tables, we have the following.
@@
    $\alpha:=\inf\left\{z\geq 1:4\varepsilon^{2}\geq\exp\left(\frac{6+12e^{4/e}+12e^{8/e}+6e^{32/e}}{z^{1/4}}\right)-1\right\}.$
    When $t=\dfrac{n^{2}q^{2}\log(\sqrt{\alpha}(n-1))}{8\pi^{2}}\left(1-\dfrac{4}{3q}\right)^{-1}$,
@@
    $t_{\operatorname{mix}}(\varepsilon)\geq\frac{n^{2}q^{2}}{8\pi^{2}}\left(\log(n)+\log\left(1-\frac{1}{n}\right)+\frac{1}{2}\log\left(\frac{\varepsilon^{-1}-1}{5}\right)\right).$
Theorem 2 implies that as $n\rightarrow\infty$ with $q\gg n^{2}$, this family of continuous-time random walks on $n\times n$ contingency tables exhibits cutoff at the mixing time $\dfrac{n^{2}q^{2}\log(n)}{8\pi^{2}}$, with a cutoff window of order $\mathcal{O}(n^{2}q^{2})$.
As mentioned in Section 1.1, we also establish cutoff for a general class of random walks on the torus ${\mathbb{Z}}_{q}^{n}$. However, since it requires detailed setup and notations, we will postpone the general result to the end of Section 6.
@@ 1.4 Background
Although there were some previous applications by practitioners, this sampling method for contingency tables via the described Markov chain was first formally discussed by Diaconis and Gangolli [7] and Diaconis and Saloff-Coste (page 373 in [11]), where the entries of the contingency tables were nonnegative integers. For $m\times n$ contingency tables over ${\mathbb{Z}}_{\geq 0}$ with $m,n$ fixed, Diaconis and Saloff-Coste showed that the mixing time of the Markov chain is of order $N^{2}$, where $N:=\sum c_{i}=\sum r_{i}$ is the sum of all the entries of the table. Hernek [15] specifically investigated the model on $2\times n$ contingency tables and proved that the mixing time is polynomial in $n$ and $N$. A modified version of the chain on $m\times n$ contingency tables was considered by Chung, Graham, and Yau [4], and they showed that the mixing time is polynomial in $m,n$, and $N$.
Through the adjacency matrix of a graph, the dynamics of the Diaconis-Gangolli random walk is closely related to the simple switching of edges, commonly used in random networks to generate random graphs with prescribed degree requirements. See, for example, Milo, Kashtan, Itzkovitz, Newman, and Alon [21], Rao, Jana, and Bandyopadhyay [25], and the survey by Wormald [26].
@@
Besides contingency tables, the cutoff phenomenon has also been widely studied. Many of the first examples of cutoffs were found in the context of card shuffling, or random walks on the symmetric group. For example, the pioneering paper by Diaconis and Shahshahani [10] established cutoff for the random transposition walk on symmetric groups. Aldous and Diaconis [1] were the first to use the term “cut-off phenomenon”. Bayer and Diaconis [2] proved the cutoff for the riffle shuffle chain, a result later known as “seven shuffles suffice”.
@@ 1.5 Outline
Section 2 contains a few lemmas that will be useful later. We prove a mixing time lower bound for more general random walks on the torus ${\mathbb{Z}}_{q}^{n}$ in Section 3, and this result can be applied to both of the specific contingency table random walks. Section 4 gives detailed proof of the upper mixing time bound for the random walk on $1\times n$ contingency tables over ${\mathbb{Z}}_{q}$. We use the same proof strategy in Section 5 to prove an upper mixing time bound for more general random walks on the torus ${\mathbb{Z}}_{q}^{n}$, and together with our lower bound from Section 3, we establish cutoff for a class of general random walks on the torus. Finally, in Section 6, we apply an analogous argument to show cutoff for the random walk on $n\times n$ contingency tables.

### v2

# Cutoff for Contingency Table and Torus Random Walks with Low Incremental Correlations

@@ Abstract
We use the correlation matrix of the generating distribution to determine the mixing time for random walks on the torus $({\mathbb{Z}}/q{\mathbb{Z}})^{n}$. We present our method in the context of the Diaconis-Gangolli random walk on both the $1\times n$ and $m\times n$ contingency tables over ${\mathbb{Z}}/q{\mathbb{Z}}$. In the $1\times n$ case, we prove that the random walk exhibits cutoff at time $\dfrac{nq^{2}\log(n)}{8\pi^{2}}$ when $q\gg n$; in the $m\times n$ case, where $m,n$ are of the same order, we establish cutoff for the random walk at time $\dfrac{mnq^{2}\log(mn)}{16\pi^{2}}$ when $q\gg n^{2}$. Our method reveals that a general class of random walks on the torus $({\mathbb{Z}}/q{\mathbb{Z}})^{n}$ has cutoff. If each coordinate of the lifted random walk onto ${\mathbb{Z}}^{n}$ has variance $\sigma^{2}/n$ in each jump, and the between-coordinate correlations are sufficiently low, then cutoff occurs at time $\dfrac{nq^{2}\log(n)}{4\pi^{2}\sigma^{2}}$.
@@ 1.1 Motivation and Basic Definitions
We are interested in showing that a class of torus random walks has cutoff primarily based on the correlation matrix of the incremental distribution of the walk. We consider a continuous-time random walk $\{X_{t}\}_{t\geq 0}$ on the torus $({\mathbb{Z}}/q{\mathbb{Z}})^{n}$ when $q$ is large with respect to $n$, and the correlations between coordinates in the incremental distribution are sufficiently small. We provide sufficient conditions for cutoff based on the correlation matrix of $X_{t}$ when lifted to $\mathbb{Z}^{n}$ and the decay of its characteristic function. This setup covers various random walks currently studied in the literature, including the Diaconis-Gangolli random walk on contingency tables over ${\mathbb{Z}}_{q}:={\mathbb{Z}}/q{\mathbb{Z}}$. We will use the Diaconis-Gangolli random walk as a primary model to present our method.
@@
A contingency table is a matrix of integer entries with prescribed row and column sums. Enumerating the contingency tables with given marginals is closely related to counting a variety of other combinatorial objects, including permutations with descent restrictions, double cosets of the symmetric groups, certain quantities about the induced representations of the symmetric groups, and some special types of Young tableaux. The name “contingency table” is commonly used in the context of statistics and is linked to Fisher’s exact test of association and goodness of fit, along with the approximate $\chi^{2}$ test. Random sampling from the set of contingency tables with fixed margins can be used to form an approximate p-value for the test of association when the $\chi^{2}$ approximation is not well met.
Markov chain Monte Carlo (MCMC) methods can be used to randomly sample and approximately count contingency tables with fixed row and column sums. One of the famous MCMC algorithms was described in [7] and [11], giving rise to the Diaconis-Gangolli random walk on contingency tables over ${\mathbb{Z}}_{q}$. Because the sums of rows and columns are fixed, we can view the dynamics as a random walk on the torus by ignoring the last row and column. We will first use the Diaconis-Gangolli random walk on $1\times n$ contingency tables over ${\mathbb{Z}}_{q}$ to outline how we can use incremental correlations to establish cutoff for torus random walks (Theorem 1). Then we will present our method in full generality to prove cutoff for a general class of random walks on the torus ${\mathbb{Z}}_{q}^{n}$ satisfying assumptions on the correlations between the coordinates of a single increment (Theorem 3). Finally, we apply our method to the Diaconis-Gangolli random walks on the more complicated $m\times n$ contingency tables over ${\mathbb{Z}}_{q}$ and illustrate the cutoff phenomenon in this model (Theorem 2).
In particular, the last part of this paper could be viewed as a complement to a previous work by Nestoridi and Nguyen [25]. They established cutoff for the discrete-time Diaconis-Gangolli random walk on $n\times n$ contingency tables over ${\mathbb{Z}}_{q}$, as $n\to\infty$. When $m=n$, our Theorem 2 gives cutoff for this model in continuous time when $q$ is sufficiently large relative to $mn=n^{2}$, while Nestoridi and Nguyen needed $\log(q)=o\left(\dfrac{\sqrt{\log(n)}}{\log\log n}\right)$ to achieve cutoff. Note that we need $q$ to grow faster than $n^{2}$ to establish cutoff, whereas Nestoridi and Nguyen required $q$ to be very small.
@@
In addition, our general result Theorem 3 expands on examples of cutoff for non-local random walks on $\mathbb{Z}_{q}^{n}$. Previously, Nestoridi [24] found mixing times and regimes of cutoff for a non-local random walk on ${\mathbb{Z}}_{2}^{n}=({\mathbb{Z}}/2{\mathbb{Z}})^{n}$ such that $k$ coordinates are randomly selected and flipped in a single iteration. In contrast, our results show cutoff for random walks on ${\mathbb{Z}}_{q}^{n}$ with non-nearest-neighbor steps when both $n\to\infty$ and $q=q(n)\to\infty$.
@@
We now define the dynamics of the Diaconis-Gangolli random walk, and note that we will study this model in continuous time. We start with the $1\times n$ contingency tables.
@@
    Flip a fair coin. If heads, increment the first coordinate selected by +1+1 and increment the second coordinate selected by 1-1; if tails, increment the first coordinate selected by 1-1 and increment the second coordinate selected by +1+1.
The analogous definition and dynamics for $m\times n$ contingency tables are the following.
@@
We will study the rate-1 continuous-time Diaconis-Gangolli random walk on $m\times n$ contingency tables over ${\mathbb{Z}}_{q}$, each of whose jumps is governed by the following rule:
@@
    Flip a fair coin, and consider the $m\times n$ matrix $T=T(i,i^{\prime},j,j^{\prime})$ whose $(i,j)$-th and $(i^{\prime},j^{\prime})$-th entries are +1+1, $(i,j^{\prime})$-th and $(i^{\prime},j)$-th entries are 1-1, and zero elsewhere,
@@
    If heads, add $T$ to the current matrix; if tails, add T-T to the current matrix.
As mentioned above, we specifically consider the case when $m$ has the same order $n$. More precisely, we assume that there are universal constants $\kappa_{1},\kappa_{2}$ such that $0<\kappa_{1}\leq 1\leq\kappa_{2}$ and
$\kappa_{1}n\leq m\leq\kappa_{2}n.$
For convenience, we will also assume that $\kappa_{1}n,\kappa_{2}n\in{\mathbb{Z}}$, and $\kappa_{1}(n-1)\leq m-1\leq\kappa_{2}(n-1)$. Since we will take $n\to\infty$, these assumptions will not affect our results.
@@
for $\varepsilon\in(0,1)$. In particular, we show that both the walks on $1\times n$ and $m\times n$ contingency tables exhibit cutoff as $n\to\infty$. We say that a sequence of Markov chains indexed by $n=1,2,\ldots$ has cutoff if for all $\varepsilon\in(0,1)$,
@@
In other words, for large enough $n$, there will be a sharp transition for $d(t)$ from almost $1$ to almost 0.
@@ 1.2 Notations
Throughout the paper, we will use ${\mathbb{Z}}_{q}$ to denote the cyclic group ${\mathbb{Z}}/q{\mathbb{Z}}$ of $q$ elements. We do not need $q$ to be a prime number, but we intrinsically require $q=q(n)$ to increase along with $n$. In the case of $m\times n$ contingency table, $m=m(n)$ should also be considered as a function of $n$ such that $\kappa_{1}n\leq m\leq\kappa_{2}n$.
@@
  • $f={\mathcal{O}}(g)$, or equivalently, $f\lesssim g$, if $\limsup f/g<\infty$, and
@@
All the limits are taken as $n\to\infty$ unless otherwise specified. Lastly, we will write ${\mathbb{N}}=\{0,1,2,3,\dots\}$ for the set of nonnegative integers.
@@ 1.3 Main Results
Let $\varepsilon\in(0,1)$ be arbitrary. For the rate-$1$ continuous-time Diaconis-Gangolli random walk on $1\times n$ contingency tables over ${\mathbb{Z}}_{q}$, we have the following.
@@
Let $\varepsilon\in(0,1)$ be arbitrary, and suppose there exist universal constants $\kappa_{1},\kappa_{2}>0$ so that $0<\kappa_{1}\leq 1\leq\kappa_{2}$ and $\kappa_{1}n\leq m\leq\kappa_{2}n$. For the rate-$1$ continuous-time Diaconis-Gangolli random walk on $m\times n$ contingency tables over ${\mathbb{Z}}_{q}$, we have the following.
@@
    $\alpha:=\inf\left\{z\geq 1:4\varepsilon^{2}\geq\exp\left(\frac{\kappa_{2}^{2}(6+12e^{4/e}+12e^{8/e}+6e^{32/e})}{\kappa_{1}z^{1/4}}\right)-1\right\}.$
    When $t=\dfrac{mnq^{2}\log(\alpha(m-1)(n-1))}{16\pi^{2}}\left(1-\dfrac{4}{3q}\right)^{-1}$,
@@
    $t_{\operatorname{mix}}(\varepsilon)\geq\frac{mnq^{2}}{16\pi^{2}}\left(\log\left((m-1)(n-1)\right)+\log\left(\frac{\varepsilon^{-1}-1}{5}\right)\right).$
Theorem 2 implies that as $n\rightarrow\infty$ with $q\gg mn$, this family of continuous-time random walks on $m\times n$ contingency tables exhibits cutoff at the mixing time $\dfrac{mnq^{2}\log(mn)}{16\pi^{2}}$, with a cutoff window of order $\mathcal{O}(mnq^{2})$.
As mentioned in Section 1.1, we also establish cutoff for a general class of random walks on the torus ${\mathbb{Z}}_{q}^{n}$. We will now give a shortened version of the generalization, and see Section 5.4 for the full statement of the theorem. In our setup, we assume that $\{X_{t}\}_{t\geq 0}^{(n)}$ is an irreducible rate-1 continuous-time random walk on the torus ${\mathbb{Z}}_{q}^{n}$ with symmetric generating distribution $\mu^{(n)}$, and let $Y^{(n)}$ be a random vector defined on $\mathbb{Z}^{n}$ with the law of $\mu^{(n)}$ lifted to $\mathbb{Z}^{n}$. We assume the coordinates of $Y^{(n)}$ have a marginal incremental variance $\dfrac{\sigma^{2}}{n}$, with $\sigma^{2}=\sigma^{2}(n)$, and $Y^{(n)}$ has correlation matrix $\Gamma_{n}$. We will also assume that $\|Y^{(n)}\|_{1}\leq r=r(n)$, the sequence $\{\Gamma_{n}\}_{n=1}^{\infty}$ satisfies the Correlation Condition, and the sequence $\{Y^{(n)}\}_{n=1}^{\infty}$ meets the Decay Condition in terms of its characteristic function (see Section 5 for the formal definitions of both conditions). In our regime, we assume that both $n\rightarrow\infty$ and $q\rightarrow\infty$. Based on this setup, we can now state a shortened version of the general theorem.
Theorem 3.
Let $\{X_{t}\}^{(n)}$ be an arbitrary family of irreducible symmetric continuous-time Markov chains on the torus $\mathbb{Z}_{q}^{n}$ with the setup above, assuming $Y^{(n)}$ has correlation matrix $\Gamma_{n}$. Let $\varepsilon\in(0,1)$ be arbitrary. Then we have the following.
  1. (a)
    Suppose that $\{Y^{(n)}\}$ meets the Decay Condition and $\{\Gamma_{n}\}_{n=1}^{\infty}$ meets the correlation condition. Then there exists $\alpha=\alpha(\varepsilon)$ so that when
    $t=\dfrac{nq^{2}\log(\alpha n)}{4\pi^{2}\sigma^{2}}\left(1-\dfrac{r^{2}}{12q}\right)^{-1}$
    Then,
    $d^{(n)}(t)\leq\varepsilon+o(1)$
    as $n\rightarrow\infty$, $q\rightarrow\infty$ with $r^{2}\ll q$.
  2. (b)
    On the other hand as $n\rightarrow\infty$, $q\rightarrow\infty$ with $\sup_{k\neq\ell}|\Gamma_{n,k\ell}|\ll\dfrac{1}{\log(n)}$, and $\dfrac{r^{2}}{q^{2}}\ll\dfrac{1}{\log(n)}$,
    $t_{\operatorname{mix}}^{(n)}(\varepsilon)\geq\dfrac{nq^{2}}{4\pi^{2}\sigma^{2}}\left(\log(n)+\log\left(\frac{\varepsilon^{-1}-1}{5}\right)\right)$
We see that if the family of Markov chains $\{X^{(n)}_{t}\}_{n=1}^{\infty}$ meets the assumptions of Theorem $3$, then $\{X^{(n)}_{t}\}_{n=1}^{\infty}$ has cutoff at the mixing time $\dfrac{nq^{2}\log(n)}{4\pi^{2}\sigma^{2}}$, with a cutoff window of order $\dfrac{nq^{2}}{\sigma^{2}}$. Here our primary regularity conditions rely on the correlations of the walk being sufficiently low, and the generating distribution of the walk having a fast enough rate of decay outside a neighborhood of 0. We will postpone a more precise description of the general result to the end of Section 5.
@@ 1.4 Background
This sampling method for contingency tables via the described Markov chain had been applied in practice, but it was first formally discussed by Diaconis and Gangolli [7] and Diaconis and Saloff-Coste (page 373 in [11]), where the entries of the contingency tables were nonnegative integers. For $m\times n$ contingency tables over $\mathbb{N}$} with $m,n$ fixed, Diaconis and Saloff-Coste showed that the mixing time of the Markov chain is of order $N^{2}$, where $N:=\sum c_{i}=\sum r_{i}$ is the sum of all the entries of the table. Hernek [16] specifically investigated the model on $2\times n$ contingency tables and proved that the mixing time is polynomial in $n$ and $N$. A modified version of the chain on $m\times n$ contingency tables was considered by Chung, Graham, and Yau [4], and they showed that the mixing time is polynomial in $m,n$, and $N$.
Through the adjacency matrix of a given graph, the dynamics of the Diaconis-Gangolli random walk is closely related to the simple switching operations of the edges. Namely, we pick two non-incident edges $i\leftrightarrow j$, $i^{\prime}\leftrightarrow j^{\prime}$ uniformly at random at each step, and then we delete the existing two edges and add the edges $i\leftrightarrow j^{\prime}$, $i^{\prime}\leftrightarrow j$. The switch chain formed by these dynamics could be used to generate configuration model random graphs or to generate random graphs with prescribed degree requirements in random networks. Recently, Tikhomirov and Youssef [27] showed that the switch chain on $d$-regular bipartite graphs on $n$ vertices has a mixing time of order $(nd)^{2}\log(nd)$, under the assumption that the degree $d\geq 3$ is at most polynomial in $n$. For references in the context of random networks, see, for example, Milo, Kashtan, Itzkovitz, Newman, and Alon [22], Rao, Jana, and Bandyopadhyay [26], and the survey by Wormald [28].
@@
In some other Markov chain models, the cutoff phenomenon has also been widely observed and studied. Many of the first examples of cutoffs were found in the context of card shuffling, or random walks on the symmetric group. For example, the pioneering paper by Diaconis and Shahshahani [10] established cutoff for the random transposition walk on symmetric groups. Aldous and Diaconis [1] were the first to use the term “cut-off phenomenon”. Bayer and Diaconis [2] proved the cutoff for the riffle shuffle chain, a result later known as “seven shuffles suffice”.
@@ 1.5 Outline
Section 2 contains a few lemmas that will be useful later. We prove a mixing time lower bound for more general random walks on the torus ${\mathbb{Z}}_{q}^{n}$ in Section 3, and this result can be applied to both of the specific contingency table random walks. Section 4 gives detailed proof of the upper mixing time bound for the random walk on $1\times n$ contingency tables over ${\mathbb{Z}}_{q}$. We use the same proof strategy in Section 5 to prove an upper mixing time bound for more general random walks on the torus ${\mathbb{Z}}_{q}^{n}$, and together with our lower bound from Section 3, we establish cutoff for a class of general random walks on the torus. Finally, in Section 6, we apply an analogous argument to show cutoff for the random walk on $m\times n$ contingency tables.